Rotational Motion Problem Solving An expert is a person who has made all the mistakes that can be made in a very narrow fielplease help with problem 53
Answer #1

a. I hope you can easily draw the freebody diagram

There are mainly 3 forces acting on the body

1. Frictional force acting parallel to the plane and in the opposite direction of motion

2. Normal force which is perpendicular to the plane

3. Weight of the body that is gravitational force

Which can be resolved into two a cos\Theta component acting in opposite direction of normal force and cancels it

And a sin\Theta component opposite to frictional force

B. Here torque = f × r ( f = frictional force' r = radius of sphere ) = force × perpendicular distance

We have \tau = Io × \alpha where Io = moment of inertia and \alpha = angular acceleration

\tau= (2MR^2 /5) × a/R = 2MRa /5

We could find a from this equation but \tau is not given so we substitute \tau by f R we get  f = 2Ma/5 equation 1

we can calculate f from equation of translational motion

According to newton's law  ma = f - mgsin\Theta ( f= \mu × normal force = \mu mgcos\theta)

  f = m(a+gsin\Theta) equation 2

from equation 1 and 2

2Ma/5 = Ma+Mgsin\Theta

a = -5gsin\Theta/3

Putting these result into equation 2

We get f= -2gsin\Theta/3

C. hollow sphere has larger moment of inertia than a solid sphere it will have smaller acceleration and will take more time than solid sphere to reach gtounf